OpenStudy (anonymous):

fan/medal can i please get a step by step solution for 3 times 4 - 5 (3 x 4 - 5)

OpenStudy (anonymous):

Is that an "X" or is that a multiplaction symbol where it says 3 x 4?

OpenStudy (anonymous):

I know it equals 7, but I need a step by step

OpenStudy (anonymous):

@Soniatonia can you answer my question please?

OpenStudy (anonymous):

um, i'll try

OpenStudy (anonymous):

I mean, I asked you if 3 x 4 is multiplication or if it's the letter "X".

OpenStudy (anonymous):

oh, it's multiplication

OpenStudy (anonymous):

Okay, first.. Simplify the following: \[3 * 4 - 5 (3 * 4 - 5)\] Then.. Multiply the 3 and 4 together inside the parenthese's.. \[3*4 = 12\] Leaving us with \[3*4-5 (12-5)\] Then subtract 5 from 12 inside the parenthese's.. \[12-5=7\] Leaving us with \[3*4 -5 (7)\] Know where to go from here?

OpenStudy (anonymous):

3∗4−5(3∗4−5) 3*4=12 12-5(12-5) 12-5(7) 12-35 -23 Answer is -23

OpenStudy (anonymous):

is that it,@Dakotafox79?

OpenStudy (anonymous):

What @DIDITHEDAD said is correct, the right answer is -23.. I was just going more in depth than what they were.

OpenStudy (anonymous):

oh, thank you both :)